OpenStudy (bekkah323):

Equation below 5x-5/16 divided by x-1/6

OpenStudy (bekkah323):

\[\frac{ 5x-5 }{ 16 }\div \frac{ x-1 }{ 6 }\]

OpenStudy (anonymous):

first get the reciprocal and multiply \[\large \frac{5x - 5}{16} \times \frac 6{x-1}\]

OpenStudy (bekkah323):

yah i got that far i just don't know how to multiply it from there @xlegendx

OpenStudy (anonymous):

don't multiply yet factor and cancel

OpenStudy (anonymous):

\[\large \frac{5x - 5}{16} \times \frac 6{x-1}=\frac{5(x-1)\times 6}{16\times (x-1)}\]

OpenStudy (anonymous):

away go the \(x-1 \) factors leaving only \(\frac{30}{16}=\frac{15}{8}\)